OpenStudy (anonymous):

What is the final velocity of a rocket propelled into space with a net acceleration of 200m/s^2 as it reaches an altitude of 10.0km? (Assuming an initial velocity of zero) ***thank you!!

OpenStudy (anonymous):

Is this kinematics, or energy?

OpenStudy (anonymous):

more kinematics:)

OpenStudy (anonymous):

First look at velocity: \[ v(t)-v(0) = \int_0^ta(x)~dx = \int_0^t200(m/s^2)~dx = 200(m/s^2)\bigg|_{x=0}^{x=t} \]Initial velocity is zero, meaning \(v(0)=0\). We end up with:\[ v(t) = 200t(m/s^2) \]

OpenStudy (anonymous):

Now we need to find \(t_f\), which is the time it reaches that high altitude.

OpenStudy (anonymous):

okay:)

OpenStudy (anonymous):

\[ x(t_f)-x(0) = \int_0^{t_f}200t(m/s^2)~dt = 200\frac{t^2}{2}(m/s^2)\bigg|_{t=0}^{t=t_f} \]We know that \(x(t_f)-x(0)=10.0(km)\) since that was its increase in altitude. So we get: \[ 10.0(km) =200\frac{(t_f)^2}2(m/s^2) \]We can use this to solve for \(t_f\).

OpenStudy (anonymous):

ohh okay, so from here, would I first multiply by 2 on both sides? 20.0 (km)=200(tf)^2 ?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

and then now, divide by 200? 0.1= tf^2 ?

OpenStudy (anonymous):

and square and get tf=0.316 ?

OpenStudy (anonymous):

Yes, but remember your units. \[ 0.1(km) = t_f^2(m/s^2) \]

OpenStudy (anonymous):

ahh right!! okay, and then tf=0.316 km/s/s^2 ?

OpenStudy (anonymous):

Hold on, what units should \(t_f\) be in?

OpenStudy (anonymous):

m/s^2 ?

OpenStudy (anonymous):

Nope.

OpenStudy (anonymous):

ohh darn :( not sure!! km?

OpenStudy (anonymous):

First divide both sides by the units. \[ \frac{0.10(km)}{(m/s^2)}=t_f^2 \]

OpenStudy (anonymous):

okay, so 0.10 km = t^f(m/s^2) ?

OpenStudy (anonymous):

Remember that \(1km = 1000m\)

OpenStudy (anonymous):

so 100 m = tf^2(m/s^2) ?

OpenStudy (anonymous):

Yeah \[ 100m = t_f^2(m/s^2) \]First divide by \(m\) to get: \[ 100 = t_f^2(1/s)^2 \]Then multiply by \(s^2\) to get: \[ 100(s^2)=(t_f)^2 \]Then you take the square root.

OpenStudy (anonymous):

ahhh okie:) and then we get tf=10 ?

OpenStudy (anonymous):

\[ t_f=10(s) \]

OpenStudy (anonymous):

ohh okay:) and so that would be the final velocity? :O

OpenStudy (anonymous):

Then we need to put that into our equation we got for velocity:\[ v(t) = 200t(m/s^2) \]

OpenStudy (anonymous):

\[ v_f=v(t_f) = 200(10(s))(m/s^2) \]

OpenStudy (anonymous):

oh! okay:) and that's the final velocity? :O plug in and done? :O

OpenStudy (anonymous):

Simplify it.

OpenStudy (anonymous):

so we get 2000 m/s ?

OpenStudy (anonymous):

@wio? did i simplify that correctly? would 2000 m/s be the final velocity?

OpenStudy (anonymous):

I think so.

OpenStudy (anonymous):

okie!! yay! thank you very much!! :D